7. A baseball bat starts from rest and is swing through 140 ° in 0.7 s. What is the angular acceleration?

A. 200 °/s2 B. 189.1 °/s2 C. 564.3 °/s2 D. 571.4 °/s2

8. A diver starts from rest and accelerates through a somersault at 950 °/s2. What is the angular displacement over 0.15 s?

A. 10.69° B. 142.5° C. 71.25 °D. 950°

9. A golf club started from rest and reached a final angular velocity of 370 °/s with a constant angular acceleration of 580°/s2. How many degrees did the club rotate?

A. 210 B.118 C156 D.213
Answer #1

Question 7

Initial angular speed of the baseball bat = = 0 o/s (At rest)

Time period = T = 0.7 sec

Angle through which the baseball bat swings = = 140 o

Angular acceleration of the baseball bat =

= T + T2/2

140 = (0)(0.7) + (0.7)2/2

= 571.4 o/s2

Angular acceleration of the baseball bat = 571.4 o/s2

Question 8

Initial angular speed of the diver = = 0 o/s (At rest)

Angular acceleration of the diver = = 950 o/s2

Time period = T = 0.15 sec

Angular displacement of the diver =

= T + T2/2

= (0)(0.15) + (950)(0.15)2/2

= 10.69 o

Angular displacement of the diver over 0.15 sec = 10.69 o

Question 9

Initial angular velocity of the golf club = 1 = 0 o/s (At rest)

Final angular velocity of the golf club = 2 = 370 o/s

Angular acceleration of the golf club = = 580 o/s2

Angle through which the golf club rotated =

22 = 12 + 2

(370)2 = (0)2 + 2(580)

= 118 o

Angle through which the club rotates = 118 o